# S4ST: A Strong, Self-transferable, faSt, and Simple Scale Transformation for Data-free Transferable Targeted Attack

### v2

@@ 3.4.2 S4ST-Base
We define the intensity of the S4ST-Base with the variable $r$, which is distinct from the definition of $s$ in Appendix A. Consider a sample $\bm{x}$ with dimensions $H\times W$ and an intensity $r>1$. With a probability $p_{r}$, S4ST-Base selects a relative scale factor $r^{\prime}$ from a uniform distribution $\mathcal{U}(\frac{1}{r},r)$. If $r^{\prime}<1$, the sample $\bm{x}$ is downscaled to dimensions $r^{\prime}H\times r^{\prime}W$, followed by random zero-padding to restore its original size. In contrast, if $r^{\prime}>1$, a square patch of dimensions $\frac{1}{r^{\prime}}H\times\frac{1}{r^{\prime}}W$ is randomly cropped from the image and then upscaled to the initial dimensions.
Scaling transformations are also utilized in existing approaches, such as DI [10] and RDI [28]. The innovations of our work compared to these methods are threefold: 1) Our black-box condition analysis singles out scaling as the foremost effective basic transformation other than empirical choice; 2) We ascertain that an extended range of scale transformations can lead to further enhancements; 3) We introduce a versatile scaling transformation that includes both enlargement and reduction, unlike previous methods that typically focus on one direction. These advancements enable our S4ST-Base to stand out for its combined effectiveness and efficiency. A comprehensive comparison among variations of the scaling transformation is presented in Section 4.2.3.
3.${}^{\textit{4}}$-Aug
@@
Guided by the insights gained from Fig. 9, we aimed to integrate a suite of complementary basic transformations into the S4ST-Base to further boost its efficacy. With a probability $p_{\textit{aug}}$, S4ST-Aug initially applies a randomly chosen transformation from an equitable selection pool of basic transformations to the entire input image. To compile this pool, we execute S4ST-Base across a spectrum of intensities $r\in[1.02,1.04,1.06,1.08,1.1,1.2,1.4,1.6,1.8,2.0,2.2]$. We then evaluate the linear correlation of self-transferability for this array of AEs over a range of basic transformations, with the results displayed in Fig. 11. We select five basic transformations that demonstrate the lowest PCCs with $\textit{ST}_{\text{Scaling}}$. Additionally, we omit the translate and crop transformations due to their significant overlap with S4ST-Base. Following the parameterization in torchvision [^1] https://github.com/pytorch/vision/tree/main/torchvision/transforms, the transformation pool for S4ST-Aug includes the following basic transformations with their empirically determined parameters:
@@
3.4.${}^{\textit{4}}$-Block
Recent research has underscored the value of block-wise transformations in enhancing adversarial transferability. With S4ST-Block, we advance this concept by randomly dividing the image into $m$ blocks and applying S4ST-Base to each block, where $m=m_{h}\times m_{w}$ indicates partitioning the image into $m_{h}$ and $m_{w}$ segments along the coordinate axes (not specifying horizontal or vertical orientation in order).
We acknowledge that the design of the block-wise transformation is heavily inspired by prior works such as SIA [11] and BSR [12]. It is important to reiterate that our design is rooted in the analysis conducted within this study, as opposed to being purely based on intuition or empirical choice, in accordance with the shortcomings of existing paradigms discussed in Section 2.4. This analytical approach yields notable differences between our proposed method and pre-existing ones: 1) Our method solely concentrates on block-wise scaling, deduced from our self-alignment estimation concerning the efficacy of various basic transformations, unlike SIA’s arbitrary transformations for each image block or BSR’s shuffling and rotation operations; 2) The selection of complementary transformations in our approach is guided by correlation analyses of self-transferability under black-box scenarios (as depicted in Fig. 11), rather than the intuitive selections employed by SIA and BSR. Comparisons against these methods are provided in Section 4.3.
@@ 3.4.5 Black-box Parameter Selection Strategy
Unlike most existing methods that rely on accessing black-box models for hyperparameters tuning, the parameter configuration of the proposed S4ST method also adheres to the strict black-box setting. First, our parameter optimization objective is to maximize the average self-transferability across the 12 studied basic transformations at different intensities, rather than black-box transferability involving the victim models. Second, we employ Bayesian optimization to simultaneously search for the four parameters of S4ST: $[p_{\textit{r}},r,p_{\textit{aug}},m]$, avoiding suboptimal results from optimizing individual components separately. The optimization is performed on a randomly selected subset of the ImageNet-Compatible dataset (50 images), consisting of 10 initial random explorations and a total of 100 trials. The final resulting setting is $p_{\textit{r}}=0.9,r=1.7,p_{\textit{aug}}=1.0,m=6$.
@@ 4 Experiments
In this section, we detail the experimental setup, and scrutinize the analysis and the ablation of components for our proposed method, showcasing its novel aspects in contrast to existing techniques. This is followed by extensive experiments designed to compare our method with the existing transformation techniques and SOTA approaches for TTAs.
@@ 4.1 Experimental Settings
Dataset. Our experiments are conducted on the ImageNet-Compatible dataset [^2] https://github.com/cleverhans-lab/cleverhans/tree/master/cleverhans_v3.1.0/examples/nips17_adversarial_competition/dataset, a benchmark commonly utilized for evaluating TTA performance [24, 33, 32]. Released for the NIPS 2017 adversarial competition, it encompasses 1,000 images with dimensions of 299$\times$299, each paired with a targeted label. For our evaluated models, we resize the images to 224$\times$224.
Models. In line with prior studies, we predominantly generate AEs using ResNet-50 [53] and transfer them to a variety of dissimilar victim models. These include CNN architectures—{MobileNet-v2 (MNv2) [54], EfficientNet-b0 (EN) [55], ConvNeXt-small (CNX) [56], Inception-v3 (INv3) [57], Inception-v4 (INv4) [58], Inception ResNet-v2 (IRv2), and Xception (Xcep) [59]} and vision transformers—{ViT [60], SwinT [61], MaxViT [62], Twins [63], PiT [64], TNT [65], and DeiT [66]}. We source their pre-trained weights from the torchvision [^3] https://github.com/pytorch/vision/tree/master/torchvision/models and timm [67] libraries. Supplemental black-box transfer experiments are performed on defense models and real-world APIs, with further details provided in subsequent sections.
Competitors. As for existing transformation techniques, we compare our S4ST with DI [10], RDI [28], SI [44], Admix [45], SSA [46], ODI [32], SIA [11], BSR [12], T-Aug [27], and H-Aug [27]. We also compare S4ST-TMI with SOTA solutions including SASD-WS-DI-TMI [21], DRA-DI-TMI [20], IAA [19], Clean Feature Mixup (CFM-RDI-TMI) [29], Self-Universality (SU-DI-TMI) [33], Little Robust Surrogate (LRS-DI-TMI) [35], Dominant Feature Attack (DFA) [14], TTP [15], M3D [18], and C-GSP [37]. We adhere to the implementations detailed in their respective publications and the TransferAttack [^4] https://github.com/Trustworthy-AI-Group/TransferAttack code base.
Metrics. We gauge the efficacy of different methods by their average tSuc, i.e., the proportion of successful transfers across the dataset from the surrogate to the victim models. Additionally, we evaluate the time required by different transformations to craft AEs. All experiments are conducted on an NVIDIA GeForce RTX 4090 GPU using the PyTorch [68] framework. For consistency, the batch size is set to 10 across all experiments, ensuring that no method surpasses the memory threshold, thereby maintaining fairness in the comparison of time consumption metrics.
[…]
@@ 4.2.1 Component Ablation
We first detail the ablation outcomes for the components of S4ST with previously reported configuration, as illustrated in Tables II and III. Remarkably, S4ST-Aug and S4ST-Block exhibit satisfying synergy and collectively contribute to a notable increase in performance, achieving an 83.0% average tSuc. Additionally, the pool for S4ST-Aug, which consists of five transformations, demonstrates a degree of complementarity and robustness, with the exclusion of any single transformation having a minimal impact on the overall performance.
TABLE II: Results by different combinations of the S4ST components.
S4ST-Base S4ST-Aug S4ST-Block $[p_{\textit{r}},r,p_{\textit{aug}},m]$ tSuc Time
@@
TABLE III: Results by excluding complementary augmentations from S4ST.
S4ST -Flip -Contrast -Brightness -Saturation -Hue
83.0 80.8 82.6 82.7 82.7 82.6
4.2.2 Scaling versus Other Basic Transformations
Here we validate the selection rationale for the scaling transformation, evaluating the judiciousness of transformation choices through the lens of the proposed self-alignment evaluation. To this end, we generate adversarial samples employing each basic transformation under consideration. With a probability of 0.9, given a transformation strength $s$, each transformation is executed with an intensity $s^{\prime}\in\mathcal{U}(0,s)$, selecting a random variant for the surrogate model to process and compute gradients. The outcomes illustrated in Fig. 12 reveal that scaling transformation stands out, achieving an unparalleled average tSuc of 58.0% at $s=0.8$. This performance is nearly 10% superior to that of the shear transformation (48.8%) and the perspective transformation (48.4%). Notably, the shear transformation shows substantial effectiveness at lower intensities; however, its efficiency markedly wanes at higher intensities. Furthermore, it is observed that employing color transformations in isolation offers marginal benefits for the attack. Yet, subsequent experimentation will demonstrate that when synergized as auxiliary augmentations to scaling, color transformations can significantly bolster targeted transferability.
4.2.3 Comparison among Scaling’s Variations
Figure 13: Our analysis uncovers significant untapped potential in existing scaling-related methods. We illustrate that prior achievements can be substantially enhanced by extending the scaling range and incorporating bidirectional scaling. Our tailored refinements to the S4ST-Base further amplify this effect.
We perform comparative evaluations of the (R)DI methods under both unidirectional (i.e., limited to either enlargement or reduction) and bidirectional conditions (i.e., integrating both enlargement and reduction), with the findings presented in Fig. 13. Specifically, under the hyperparameter configurations from prior TTA research, the DI and RDI methods attain an average tSuc of only 22.4% and 27.3%, respectively. Our research reveals that minor adjustments to their parameters can elevate these figures to 36.6% and 41.7%. By synergistically applying these methods in both the upscale and downscale directions, we obtain a significantly enhanced average tSuc of 56.8% at $r=2.4$, which, while slightly lower than the 57.2% achieved by the simple scaling transformation, is substantially surpassed by our S4ST-Base approach, which further boosts the outcome to 61.6%. This highlights the superiority of our method in both unlocking and augmenting the effectiveness of the basic scaling transformation. Fig. 14 delineates the time consumption of these methods, highlighting our approach’s computational efficiency. The incremental time investment required by S4ST-Base compared to the baseline method (at $r=1.0$) is minimal, and it markedly surpasses the competitors.
Figure 14: In addition to its effectiveness, our S4ST also exhibits superior efficiency. The additional time required by our S4ST is trivial, significantly outstripping the simple scaling and the integration of RDI [28] and DI [10] methods in terms of computational speed.
TABLE IV: Comparison between input transformations based on ResNet-50 as the surrogate model. strands for TTA-oriented methods. The best and the second best results are in bold and underlined, respectively.
Attack Venue tSuc Time MNv2 EN CNX INv3 INv4 IRv2 Xcep AvgC ViT Swin MaxViT Twins PiT TNT DeiT AvgT
TMI [8, 9] CVPR 2019 0.8 1.24 2.0 1.5 2.4 0.6 0.7 0.7 0.9 1.3 0.2 0.9 0.4 0.1 0.1 0.2 0.2 0.3
+Admix [45] ICCV 2021 2.1 6.27 5.9 4.7 6.3 0.9 1.6 2.4 2.5 3.5 0.3 2.0 1.0 0.5 1.1 0.6 0.2 0.8
+SI [44] ICLR 2020 4.4 6.11 8.5 9.6 8.4 5.7 5.3 6.3 7.0 7.3 1.1 2.4 1.5 1.2 2.5 1.3 0.6 1.5
+SSA [46] ECCV 2022 19.8 1.43 35.0 32.0 35.1 22.5 30.4 35.1 33.0 31.9 4.8 11.1 7.1 7.4 10.8 7.2 6.0 7.8
+DI [10] CVPR 2019 22.4 1.44 29.6 33.7 48.8 15.4 35.4 38.1 28.3 32.8 5.9 17.1 10.8 11.1 19.1 11.6 8.3 12.0
+RDI [28] ECCV 2020 27.3 1.24 37.1 42.9 47.4 25.8 39.8 45.6 36.9 39.4 11.0 21.7 13.2 13.6 22.3 15.0 10.2 15.3
+ODI [32] CVPR 2022 58.1 5.11 73.2 77.1 75.3 64.3 75.7 77.0 73.2 73.7 37.6 54.1 37.9 38.4 54.5 43.0 31.7 42.5
+SIA [11] ICCV 2023 58.3 1.55 82.5 81.3 84.3 51.1 74.2 76.2 72.0 74.5 30.8 58.5 42.6 39.3 53.2 40.7 29.8 42.1
+DeCoWA [13] AAAI 2024 58.4 2.78 74.9 78.0 77.4 56.7 74.4 77.7 73.7 73.3 40.6 58.4 36.1 36.6 55.1 44.1 33.3 43.5
+BSR [12] CVPR 2024 60.0 1.84 81.8 80.3 81.4 62.6 79.3 78.1 76.6 77.2 32.2 56.8 38.4 39.6 55.7 44.2 33.5 42.9
+T-Aug [27] ArXiv 2023 65.5 1.33 82.1 82.7 81.9 72.4 84.1 84.4 82.8 81.5 41.7 61.9 47.5 45.2 64.2 50.2 36.4 49.6
+H-Aug [27] ArXiv 2023 68.8 5.64 81.8 82.4 85.9 69.7 86.0 82.8 84.7 81.9 43.1 73.4 56.7 59.0 63.0 54.1 40.4 55.7
+S4ST Ours 83.0 1.45 93.5 93.9 93.9 85.9 93.0 92.9 93.1 92.3 63.0 86.3 74.8 73.8 80.1 75.1 63.2 73.8
@@ 4.2.3 Comparison among Scaling’s Variations
Building upon the S4ST-Base framework, we delve into the benefits conferred by the proposed complementary augmentations (S4ST-Aug) and block-wise scaling (S4ST-Block). We maintain the probability of scaling transformations at $p_{\textit{r}}=0.9$ and progressively increase the number of image blocks $m$ and the probability of applying complementary augmentations $p_{\textit{Aug}}$. The outcomes are depicted in Fig. 15. For S4ST-Block, an escalation in the number of image blocks correlates with a marked enhancement in attack efficacy, attaining its zenith at $m=6\,(2\times 3)$. A further increase in blocks does not yield improved convergence within the configured iterations for the attack. In the case of S4ST-Aug, a higher frequency of applying complementary augmentations is linked with a steady amplification in performance. Additionally, the integration of five distinct types of complementary transformations, particularly when paired with scaling (for $r\geq 1.2$), results in substantially improved outcomes compared to their solitary application (at $r=1.0$). This emphasizes the soundness of our approach, which synergizes a primary transformation with complementary enhancements.

### v3

@@ 3.4.1 S4ST-Base
Our preliminary analysis reveals scaling’s robust self-alignment across broad intensities. To unpack the mechanistic impact of scaling directionality, Fig. 14 provides a quantitative case study. We observe distinct transitions in predictive behavior across scales: large scales diminish the primary class’s dominance, allowing multiple latent categories to emerge with competing confidence scores; small scales induce highly volatile confidence distributions, reflecting the scale sensitivity that local objects dominate the global semantics. Both regimes effectively exploit the image’s inherent multi-object/multi-class information.
Figure 14: Model focus varies significantly across scales. Ground truth: vase.
Driven by these insights, S4ST-Base acts as a unified bidirectional scaling operation synergizing DI [90] and RDI [105]. Crucially, instead of naive DI upscaling, we randomly crop a local patch and resize it back to original dimensions. This mitigates the severe computational overhead of large-scale inputs while amplifying local semantic exploitation, thereby boosting adversarial transferability.
In specific, we define the intensity of S4ST-Base by $r$. Consider a sample $\bm{x}$ with dimensions $H\times W$ and an intensity $r>1$. With a probability $p_{r}$, S4ST-Base selects a relative scale factor $r^{\prime}$ from a uniform distribution $\mathcal{U}(\frac{1}{r},r)$. If $r^{\prime}<1$, the sample $\bm{x}$ is downscaled to dimensions $r^{\prime}H\times r^{\prime}W$, followed by random zero-padding to restore its original size. In contrast, if $r^{\prime}>1$, a square patch of dimensions $\frac{1}{r^{\prime}}H\times\frac{1}{r^{\prime}}W$ is randomly cropped from the image and then upscaled to the initial dimensions.
3.4.2 $^{\textit{4}}S$T-Aug
@@
S4ST-Aug integrates complementary basic transformations to boost efficacy. With probability $p_{\textit{aug}}$, it applies a randomly chosen transformation from a curated pool to the entire image. To compile this pool, we evaluate S4ST-Base AEs across intensities $r\in[1.02,1.04,1.06,1.08,1.1,1.2,1.4,1.6,1.8,2.0,2.2]$, assessing their self-transferability correlations against various basic transformations (Fig. 15). We select five transformations exhibiting the lowest PCCs with $\textit{ST}_{\text{Scaling}}$, deliberately omitting translation and cropping due to their high overlap with S4ST-Base. Following torchvision [^1] https://github.com/pytorch/vision/tree/main/torchvision/transforms parameterization, the S4ST-Aug pool includes:
@@
3.4.3 $^{\textit{4}}S$T-Block
Block-wise transformations [78, 76, 36] enhance transferability by increasing transformation diversity (Fig. 6(a)). Validating its favorable correlation with self-transferability (SIA-legend in Fig. 11), we advance this concept via S4ST-Block.
It randomly divides the image into a non-overlapping $m_{h}\times m_{w}$ grid, ensuring a minimum block size for effectiveness. The dimension-preserving S4ST-Base is then independently applied to each block, diversifying the attack while maintaining coherent final adversarial examples.
Note that block-wise application serves as an auxiliary technique rather than a conventional transformation. Its efficacy likely stems from the model’s inherent tolerance to local transformations, driven by data distribution characteristics. Relevant analysis can be found in Section 5.
@@ 3.4.4 Black-box Parameter Selection Strategy
Strictly adhering to black-box constraints, S4ST tunes hyperparameters without victim model access. First, our optimization objective maximizes average self-transferability across the 12 basic transformations at varying intensities, substituting black-box transferability. Second, we employ Bayesian Optimization (BO) to jointly search all four parameters $p_{\textit{r}},r,p_{\textit{aug}},m]$, avoiding sub-optimalities from isolated component tuning. Evaluated on a 50-image subset of ImageNet-Compatible, the 100-trial BO process (including 10 initial random explorations) takes $\sim$2 hours and 15 minutes on a single NVIDIA GeForce RTX 4090 GPU ($\sim$80 seconds/trial). The resulting configuration is $p_{\textit{r}}=0.9,r=1.7,p_{\textit{aug}}=1.0,m=6$.
Fig. 16 depicts self-transferability tracking the true black-box tSuc (across 14 models) during the search. As a surrogate objective, self-transferability effectively captures critical fluctuations, trends, and relative magnitudes of true transferability, successfully guiding the BO toward high-quality parameters.
Usage and Boundary Cases. S4ST’s hyperparameter tuning is domain-specific yet dataset- and task-agnostic. The parameters optimized here can be directly applied to multiple natural image datasets and tasks, and even extend to chest X-ray classification. However, because S4ST relies on scaling and transformations derived from natural image analysis, severe domain shifts, such as face recognition, which involves fundamentally different data distributions and learning objectives, pose challenges (detailed in Section 4.6.2).
Figure 16: Self-transferability as a blind estimator for S4ST parameter search in BO. Average tSuc is calculated over 14 black-box models with $T=300$.
3.4.5 Relation to Existing Methods
While sharing implementation-level similarities with existing methods, S4ST differs fundamentally in its analytical foundation and design rationale.
Scaling-centric Design. While DI [90] and RDI [105] empirically utilize scaling, S4ST offers a principled, bidirectional mechanism. First, our self-alignment analysis analytically justifies scaling’s dominance over other transformations. Second, unlike unidirectional approaches, S4ST-Base leverages both scale enlargement (triggering fine-grained class competition) and reduction (exploiting global semantic sensitivity). Finally, replacing naive upscaling with a random crop-and-upscale strategy eliminates quadratic overhead while enforcing localized feature exploitation, yielding a superior effectiveness-efficiency balance over existing variants (Section 4.2.3).
Complementary Augmentation and Block Processing. Though inspired by composite methods like SIA [78] and BSR [76], our design eschews empirical heuristics for principled analysis (Section 2.3). Specifically, S4ST-Aug’s complementary transformations are objectively selected via black-box self-transferability correlations (Fig. 15), favoring color adjustments over intuitive geometric choices. Additionally, S4ST-Block exclusively employs self-alignment-verified block-wise scaling, contrasting with SIA’s arbitrary transformations or BSR’s empirical rotation (see Section 4.3 for detailed comparisons).
@@ 4 Experiments
In this section, we detail the experimental setup, and scrutinize the analysis and the ablation of components for our proposed method, showcasing its novel aspects in contrast to existing techniques. This is followed by extensive experiments designed to compare our method with the existing transformation techniques and SoTA approaches, covering both in-domain and cross-domain scenarios.
@@ 4.1 Experimental Settings
Dataset. Our experiments are conducted on the ImageNet-Compatible dataset [^2] https://github.com/cleverhans-lab/cleverhans/tree/master/cleverhans_v3.1.0/examples/nips17_adversarial_competition/dataset, a benchmark commonly utilized for evaluating TTA performance [100, 82, 2]. Released for the NIPS 2017 adversarial competition, it encompasses 1,000 images with dimensions of $\times299$, each paired with a targeted label.
Models. In line with prior studies, we generate AEs using ResNet-50 [27] and transfer them to a variety of dissimilar victim models. These include CNN architectures {MobileNet-v2 (MNv2) [57], EfficientNet-b0 (EN) [68], ConvNeXt-small (CNX) [46], Inception-v3 (INv3) [66], Inception-v4 (INv4) [65], Inception ResNet-v2 (IRv2), and Xception (Xcep) [13]} and Vision Transformers (ViTs) {ViT [18], SwinT [45], MaxViT [74], Twins [14], PiT [29], TNT [24], and DeiT [72]}. We source their pre-trained weights from the torchvision [^3] https://github.com/pytorch/vision/tree/master/torchvision/models and timm [86] libraries. Supplemental black-box transfer experiments are performed on defense models, real-world APIs, VLMs, dense prediction models, with further details provided in subsequent sections.
Metrics. We measure the efficacy of different methods by their average tSuc, i.e., the proportion of successful transfers across the dataset from the surrogate to the victim models. We also measure the average $\ell_{2}$ norm of generated perturbations to avoid unfair comparison [9]. Additionally, we evaluate the time required by different transformations to craft AEs. All experiments are conducted on an NVIDIA GeForce RTX 4090 GPU using the PyTorch [53] framework. For consistency, the batch size is set to 10 across all experiments, ensuring that no method surpasses the memory threshold, thereby maintaining fairness in the comparison of time consumption metrics.
4.2 Analysis on $^{\textbf{4}}S$T
@@ 4.2.1 Component Ablation
Table II details the component ablation. S4ST-Aug and S4ST-Block exhibit strong synergy, collectively driving the average tSuc to a notable 83.0%. Furthermore, S4ST-Aug’s five-transformation pool demonstrates robust complementarity; excluding any single transformation yields minimal performance drop.
TABLE II: Component ablation of S4ST.
(a) Results by different S4ST combinations
  S4ST-Base -Aug -Block $p_{\textit{r}},r,p_{\textit{aug}},m]$ tSuc Time
$[0.9,1.7,0.,1]$ 58.2 1.24
$[0.9,1.7,1.,1]$ 76.8 1.40
$[0.9,1.7,0.,6]$ 74.3 1.31
$[0.9,1.7,1.,6]$ 83.0 1.45
(b) Results by excluding complementary augmentations from S4ST
  S4ST -Flip -Contrast -Brightness -Saturation -Hue
83.0 80.8 82.6 82.7 82.7 82.6
Figure 18: Our analysis uncovers significant untapped potential in existing scaling-related methods. We illustrate that prior achievements can be substantially enhanced by extending the scaling range and incorporating bidirectional scaling. Tailored refinements to the S4ST-Base further amplify this effect. The additional time required by our S4ST is trivial, significantly outstripping the simple scaling and the integration of RDI [105] and DI [90] methods in terms of computational speed.
4.2.2 Scaling versus Other Basic Transformations
To validate simple scaling’s superiority objectively, we evaluate AEs generated using studied basic transformations. Executing each with probability 0.9 and random intensity $s^{\prime}\sim\mathcal{U}(0,s)$, Fig. 17 shows scaling uniquely achieves an unparalleled 58.0% average tSuc at $s=0.8$, outperforming shear (48.8%) and perspective (48.4%) by nearly 10%. Notably, shear’s efficacy wanes significantly at higher intensities.
4.2.3 Comparison among Scaling Variations
Fig. 18 compares (R)DI methods under unidirectional and bidirectional settings. Under prior TTA configurations, DI and RDI attain merely 22.4% and 27.3% average tSuc, respectively. Minor parameter tuning elevates these to 36.6% and 41.7%. Synergizing both upscale and downscale directions yields 56.8% at $r=2.4$. While comparable to simple scaling (57.2%), our S4ST-Base fundamentally outperforms both, boosting tSuc to 61.6%. This underscores our method’s superiority in unlocking and augmenting basic scaling. Furthermore, Fig. 18(b) confirms S4ST-Base’s computational efficiency: its incremental time investment over the baseline ($r=1.0$) is minimal, significantly outpacing rival integration methods.
4.2.4 Effectiveness of $^{\textit{4}}S$T Components
Fig. 19 illustrates the benefits of S4ST-Aug and S4ST-Block built upon S4ST-Base (fixing $p_{\textit{r}}=0.9$). For S4ST-Block, increasing the grid blocks $m$ markedly enhances efficacy, peaking at $m=6\,(2\times 3)$; further increases impede optimization convergence within 900 iterations. For S4ST-Aug, higher application frequencies ($p_{\textit{Aug}}$) steadily amplify performance. Crucially, integrating the five complementary transformations with scaling ($r\geq 1.2$) yields substantially better outcomes than their solitary application ($r=1.0$), reinforcing the soundness of synergizing a primary transformation with complementary enhancements.
TABLE III: Comparison between input transformations based on ResNet-50 as the surrogate model. strands for TTA-oriented methods. The best and the second best results are in Red and Blue, respectively.
@@ 4.2.4 Effectiveness of $^{\textit{4}}S$T Components
[…]
TABLE IV: Results by alternative surrogate models. strands for TTA-oriented methods
[…]
TMI [16, 17] 1.7 0.4 1.86 1.5 0.2 1.66 0.0 0.0 3.99
+ODI [2] 61.2 31.3 5.57 51.0 25.7 5.26 20.0 43.9 7.70
+SIA [78] 65.2 34.5 2.39 44.2 23.2 2.30 7.9 38.8 4.63
+DeCoWA [41] 56.8 30.5 3.89 59.7 33.3 3.66 6.4 24.1 8.30
+BSR [76] 63.8 32.0 2.03 43.8 20.3 1.95 15.5 41.7 4.27
+T-Aug [83] 65.0 33.7 2.01 61.6 34.7 1.75 15.0 41.4 4.06
+H-Aug [83] 72.8 45.8 4.51 60.7 36.0 7.00 - - -
+S4ST (Ours) 86.7 62.6 2.15 76.8 54.4 1.83 50.3 81.6 4.05